OpenStudy (anonymous):

Solve: ln(x^2-2)=ln(-x)

OpenStudy (anonymous):

x^2 - 2 = - x x^2 + x - 2 = 0 (x+2) (x-1) = 0 x = -2, 1

OpenStudy (mertsj):

But since the argument cannot be negative. 1 must be discarded